OpenStudy (anonymous):

What is the value of x? 3 • (–4) • x = 0 x =

Directrix (directrix):

3 • (–4) • x = 0 -12x = 0 Multiply both sides by the reciprocal of -12 which is (-1/12) (-1/12 )* (-12x) = (-1/12) * 0 x = ? @candyswirl